Listed here are four series of separate costs measured at various volume levels Volume (Units) Series 1 Series 2 Series 3 SerAstro Co. sold 20,100 units of Its only product and incurred a $63,560 loss (ignoring taxes) for the current year as shown he3. Prepare a forecasted contribution margin income statement for 2018 that shows the expected results with the machine instal4. Compute the sales level required in both dollars and units to earn $210,000 of target pretax income in 2018 with the machiRequired Information The following information applies to the questions displayed below.j Henna Co. produces and sells two prProduct T Contribution Margin Ratio Contribution Margin Ratio Choose Numerator: Choose Denominator: Contribution margin ratio2. Assume that the company expects sales of each product to decline to 28,000 units next year with no change in unit selling3. Assume that the company expects sales of each product to Increase to 59,000 units next year with no change In unit selling
Answer #1

Please give positive ratings so I can keep answering. As per HOMEWORKLIB POLICY I have solved first 5 questions. Thanks!

Volume Series 1 Series 2 Series 3 Series 4
0 0 450            800 100
100                     800 450            800 105
200                 1,600 450            800 120
300                 2,400 450        1,600 145
400                 3,200 450        1,600 190
500                 4,000 450        2,400 250
600                 4,800 450        2,400 320
Series 1 is changing in exact proportion with volume. So it is a variable cost.
Series 2 is constant irrespective of volume. So it is a Fixed cost.
Series 3 is changing after increase in volume by every 200 units. It is a Step-wise cost. A step cost is a cost that does not change steadily with changes in activity volume, but rather at discrete points.
A step cost is a fixed cost within certain boundaries, outside of which it will change
Series 4 is changing at an uneven and irregular rate. There is no way a relation can be established between Series 4 cost and volume. So this is a Curvilinear cost. A curvilinear cost, is an expense that increases at an inconsistent rate
as production volume increases. This is an irregular cost that increases at different rates as total output increases.
Astro Co.
Ans 1 2017 Ans 2 2018 Ans 3 2018 Ans 4 2018
Units sold         20,100.00 A Variable costs per unit             16.92 K=J*60% Units sold     20,100.00 A Target Income       210,000.00 S
Sales       755,760.00 B Sell price per unit             37.60 H Sell price per unit             37.60 H Total Fixed costs       403,500.00 N
Variable costs       566,820.00 C Contribution per unit             20.68 L=H-K Variable costs per unit             16.92 K Target Contribution       613,500.00 T=S+N
Contribution Margin       188,940.00 D=B-C Fixed costs 252,500.00 F Contribution per unit             20.68 L Contribution per unit                 20.68 L
Contribution per unit                   9.40 E=D/A Additional 151,000.00 M Contribution Margin 415,668.00 Q=L*A Sales Units         29,666.34 U=T/L
Fixed costs       252,500.00 F Total Fixed costs 403,500.00 N=F+M Total Fixed costs 403,500.00 N Sell price per unit                 37.60 H
Breakeven unit         26,861.70 G=F/E Breakeven unit      19,511.61 O=N/L Net Profit     12,168.00 R=Q-N Sales Value 1,115,454.55 V=U*H
Sell price per unit                 37.60 H=B/A Breakeven value 733,636.36 P=O*H
Breakeven value 1,010,000.00 I=G*H
Variable costs per unit                 28.20 J=C/A